Express the following as a linear combination of

p1 = 3 + x + 2x2, p2 = 5 - x + 5x2, and p3 = 2 + 5x + 3x2.
1
Expert's answer
2020-04-21T14:00:46-0400

"p_1=2x^2+x+3"

"p_2=5x^2-x+5"

"p_3=3x^2+5x+2"

"x^2+x+1=a_1(2x^2+x+3)+b_1(5x^2-x+5)+c_1(2+5x+3x^2)"

"x^2(2a_1+5b_1+3c_1)+x(a_1-b_1+5c_1)+(3a_1+5b_1+2c_1)=x^2+x+1"

"2a_1+5b_1+3c_1=1;a_1-b_1+5c_1=1;3a_1+5b_1+2c_1=1"

"(3a_1+5b_1+2c_1)-(3a_1+4b_1+8c_1)=-1"

"b_1=6c_1-1" ---(1)

"2(b_1-5c_1+1)+5b_1+3c_1=1"

"7b_1-6c_1+1=0" ---(2)

Solving (1) and (2) we get

"c_1=6\/35"

Substituting this in equation(1) we get

"b_1=1\/35;a_1=6\/35"

So,

"x^2+x+1=(6\/35)(2x^2+x+3)+(1\/35)(5x^2-x+5)+(6\/35)(2+5x+3x^2)"